# A Bias–Variance Decomposition for Ensembles over Multiple Synthetic Datasets

### 06 Feb 2024 v1

@@ 3 Experiments
Figure 1: Comparison of synthetic data generation algorithms for several prediction algorithms on the California housing dataset, with 1 to 10 synthetic datasets. DDPM and synthpop achieve smaller MSE in the downstream predictions, so they were selected for further experiments. SP-P and SP-IP are the proper and improper variants of synthpop, and DDPM-KL is DDPM with KL divergence loss. 1-NN and 5-NN are nearest neighbours with 1 and 5 neighbours. The dashed black lines show the performance of each prediction algorithm on the real data, and the solid black line shows the performance of the best predictor, random forest, on the real data. The results are averaged over 3 repeats, with different train-test splits. The error bars are 95% confidence intervals formed by bootstrapping over the repeats. Linear regression was omitted, as it had nearly identical results as ridge regression. Table 3 in the Appendix contains the numbers in the plots, including ridge regression.
@@
The downstream algorithms we consider are nearest neighbours with 1 or 5 neighbours (1-NN and 5-NN), decision tree (DT), random forest (RF), gradient boosted trees (GB), a multilayer perceptron (MLP) and a support vector machine (SVM) for both classification and regression. We also use linear regression (LR) and ridge regression (RR) on regression tasks, and logistic regression (LogR) on classification tasks, though we omit linear regression from the main text plots, as its results are nearly identical to ridge regression. Decision trees and 1-NN have a very high variance, as both interpolate the training data with the hyperparameters we use. Linear, ridge, and logistic regression have fairly small variance in contrast. Appendix C contains more details on the experimental setup, including details on the datasets, and downstream algorithm hyperparameters.
3.1 Synthetic Data Generation Algorithms
First, we compare several synthetic data generation algorithms to see which algorithms are most interesting for subsequent experiments. We use the California housing dataset, where the downstream task is regression. The algorithms we compare are DDPM (Kotelnikov et al., 2023), TVAE (Xu et al., 2019), CTGAN (Xu et al., 2019) and synthpop (Nowok et al., 2016). DDPM, TVAE and CTGAN are a diffusion model, a variational autoencoder and a GAN that are designed for tabular data. We use the implementations from the synthcity library [^5] https://github.com/vanderschaarlab/synthcity for these. Synthpop generates synthetic data by sampling one column from the real data, and generating the other columns by sequentially training a predictive model on the real data, and predicting the next column from the already generated ones. We use the implementation from the authors (Nowok et al., 2016).
(a) Abalone
(b) ACS 2018
(c) California Housing
(d) Insurance
@@ 3.1 Synthetic Data Generation Algorithms
Figure 2: MSE of the ensemble of downstream predictors on the regression datasets, with varying number of synthetic datasets $m$ from either DDPM or synthpop (SP-P). Increasing the number of synthetic datasets generally decreases MSE, especially for decision trees and 1-NN. The predictors are nearest neighbours with 1 or 5 neighbours (1-NN and 5-NN), decision tree (DT), random forest (RF), a multilayer perceptron (MLP), gradient boosted trees (GB), a support vector machine (SVM) and ridge regression (RR). The black line is the MSE of the best predictor on real data. The results are averaged over 3 repeats. The error bars are 95% confidence intervals formed by bootstrapping over the repeats. We omitted linear regression from the plot, as it had almost identical results to ridge regression. Tables 4 to 7 in the Appendix contain the numbers from the plots, including ridge regression.
The results of the comparison are shown in Figure 1. We see that synthpop and DDPM with MSE loss generally outperform the other generation algorithms, so we select them for the subsequent experiments. There is very little difference between the two variants of synthpop, so we choose the “proper” variant due to its connection with the Bayesian reasoning for using multiple synthetic datasets.
3.2 Main Experiment
@@
As our main experiment, we evaluate the performance of the synthetic data ensemble on 7 datasets, containing 4 regression and 3 classification tasks. See Appendix C.1 for details on the datasets. We use both DDPM and synthpop as the synthetic data generator, and generate 32 synthetic datasets, of which between 1 and 32 are used to train the ensemble. The results are averaged over 3 runs with different train-test splits.
@@
On the regression datasets, our error metric is MSE, which is the subject of Theorem 2.2. The results in Figure 2 show that a larger number of synthetic datasets generally decreases MSE. The decrease is especially clear with synthpop and downstream algorithms that have a high variance like decision trees and 1-NN. Low-variance algorithms like ridge regression have very little if any decrease from multiple synthetic datasets. This is consistent with Theorem 2.2, where the number of synthetic datasets only affects the variance-related terms.
@@
On the classification datasets, we consider 4 error metrics. Brier score (Brier, 1950) is MSE of the class probability predictions, so Theorem 2.2 applies to it. Cross entropy is a Bregman divergence, so Theorem B.2 from Appendix B applies to it. We also included accuracy and area under the ROC curve (AUC) even though our theory does not apply to them, as they are common and interpretable error metrics, so it is interesting to see how multiple synthetic datasets affect them. We use their complements in the plots, so that lower is better for all plotted metrics. We only present the Brier score results in the main text in Figure 3, and defer the rest to Figures 5, 6 and 7 in Appendix D.
(a) Adult
(b) Breast Cancer
(c) German Credit
Figure 3: Brier score of the ensemble of downstream predictors with varying numbers of synthetic datasets $m$, generated by either DDPM and synthpop, and averaging either probabilities or log probabilities, on the classification datasets. Increasing the number of synthetic datasets generally decreases the score, except in cases where one synthetic dataset is already enough to meet the real data performance. Probability averaging is generally outperforming log probability averaging. The predictors are logistic regression nearest neighbours with 1 or 5 neighbours (1-NN and 5-NN), decision tree (DT), random forest (RF), a multilayer perceptron (MLP), gradient boosted trees (GB), a support vector machine (SVM) and logistic regression (LogR). The black line show the loss of the best downstream predictor trained on real data. The results are averaged over 3 repeats with different train-test splits. The error bars are 95% confidence intervals formed by bootstrapping over the repeats. Tables 8 to 10 in the Appendix contain the numbers from the plots.
@@
Because Theorem B.2 only applies to cross entropy when averaging log probabilities instead of probabilities, we compare both ways of averaging.
@@
The results on the classification datasets in Figure 3 are similar to the regression experiment. A larger number of synthetic datasets generally decreases the score, especially for the high-variance models. Logistic regression is an exception, as it has low variance, as are algorithms that match the performance on the real data already with one synthetic dataset.
Probability averaging generally outperforms log probability averaging on Brier score and cross entropy. There is very little difference between the two on accuracy and AUC, which suggests that probability averaging is generally better than log-probability averaging.
@@
In Appendix D.1, we estimate the MV and SDV terms of the decomposition in Theorem 2.2. These confirm our claims on the model variances: decision trees and 1-NN have a high variance, while linear, ridge and logistic regression have a low variance.
3.3 Predicting MSE from Two Synthetic Datasets
@@
Next, we evaluate the predictions our rule of thumb from Section 2.3 makes. To recap, our rule of thumb predicts that the maximal benefit from multiple synthetic datasets is $2(\mathrm{MSE}_{1}-\mathrm{MSE}_{2})$, and $m$ synthetic datasets achieve a $1-\frac{1}{m}$ fraction of this benefit.
@@
Figure 4 contains the results for the ACS 2018 dataset, and Figures 8 to 11 in the Appendix contain the results for the other datasets. The rule of thumb is very accurate on ACS 2018 and reasonably accurate on the other datasets. The variance of the prediction depends heavily on the variance of the errors computed from the test data.
We also evaluated the rule on random forests without synthetic data, as it also applies to them. In this setting, the number of trees in the random forest is analogous to the number of synthetic datasets. We use the same datasets as in the previous experiments and use the same train-test splits of the real data. The results are in Figure 12 in the Appendix. The prediction is accurate when the test error is accurate, but can have high variance.
Figure 4: MSE prediction on the ACS 2018 dataset. The prediction is very accurate on this dataset. The solid lines for DDPM and synthpop (SP-P) show the same MSEs as Figure 2, while the dashed lines show predicted MSEs. 1-NN and 5-NN are nearest neighbours with 1 or 5 neighbours. We omitted downstream algorithms with uninteresting flat curves. See Figure 8 in the appendix for the full plot, and Figure 9 plots for the other datasets. Table 11 in the Appendix contains the numbers from the plots. The results are averaged over 3 repeats with different train-test splits. The error bands are 95% confidence intervals formed by bootstrapping over the repeats.
@@ 4 Discussion
Our theory assumes that the synthetic datasets are generated i.i.d. given either the real data or a DP summary. This means our theory will not apply to more complicated schemes for generating multiple synthetic datasets, like excluding generators based on some criteria (Dikici et al., 2020) or collecting the generators from different points of a single training run (Wang et al., 2016). However, i.i.d. synthetic datasets are still very relevant, as they are the simplest case, and uncertainty estimation techniques for statistical inference with synthetic data (Raghunathan et al., 2003; Räisä et al., 2023a) require i.i.d. synthetic datasets.
@@ Conclusion
We derived bias-variance decompositions for using synthetic data in several cases: for MSE or Brier score with i.i.d. synthetic datasets given the real data and MSE with i.i.d.  synthetic datasets given a DP summary of the real data. The decompositions make actionable predictions, such as yielding a simple rule of thumb that can be used to select the number of synthetic datasets. We empirically examined the performance of generative ensembles on several real datasets and downstream predictors, and found that the predictions of the theory generally hold in practice. These findings significantly increase the theoretical understanding of generative ensembles, which is very limited in prior literature.
@@ Appendix A Missing Proofs
$\operatorname*{\mathbb{E}}_{D_{r},D_{s}^{1:m}}[\hat{g}(x;D_{s}^{1:m})]=\frac{1}{m}\operatorname*{\mathbb{E}}_{D_{r}}\operatorname*{\mathbb{E}}_{D_{s}^{1:m}|D_{r}}\left[\sum_{i=1}^{m}g(x,D_{s}^{i})\right]=\operatorname*{\mathbb{E}}_{D_{r}}\operatorname*{\mathbb{E}}_{D_{s}|D_{r}}[g(x,D_{s})]=\operatorname*{\mathbb{E}}_{D_{r},D_{s}}[g(x,D_{s})],$ (33)
@@
Appendix B Bias-Variance Decomposition for Bregman Divergences
B.1 Background: Bregman Divergences

### 23 May 2024 v2

@@ 3 Experiments
The downstream algorithms we consider are nearest neighbours with 1 or 5 neighbours (1-NN and 5-NN), decision tree (DT), random forest (RF), gradient boosted trees (GB), a multilayer perceptron (MLP) and a support vector machine (SVM) for both classification and regression. We also use linear regression (LR) and ridge regression (RR) on regression tasks, and logistic regression (LogR) on classification tasks, though we omit linear regression from the main text plots, as its results are nearly identical to ridge regression. Decision trees and 1-NN have a very high variance, as both interpolate the training data with the hyperparameters we use. Linear, ridge, and logistic regression have fairly small variance in contrast. Appendix D contains more details on the experimental setup, including details on the datasets, and downstream algorithm hyperparameters.
3.1 Effect of Multiple Synthetic Datasets
(a) Regression datasets
(b) Classification datasets
@@ 3.1 Effect of Multiple Synthetic Datasets
Figure 1: MSE on regression datasets (a) or Brier score on classification datasets (b) of the ensemble of downstream predictors, with varying number of synthetic datasets $m$ from synthpop. Increasing the number of synthetic datasets generally decreases both metrics, especially for decision trees and 1-NN. The predictors are nearest neighbours with 1 or 5 neighbours (1-NN and 5-NN), decision tree (DT), random forest (RF), a multilayer perceptron (MLP), gradient boosted trees (GB), a support vector machine (SVM), ridge regression (RR) and logistic regression (LogR). The black line is the MSE of the best predictor on real data. The results are averaged over 3 repeats. The error bars are 95% confidence intervals formed by bootstrapping over the repeats. Tables S4 to S10 in the Appendix contain the numbers from the plots.
@@
As our first experiment, we evaluate the performance of the synthetic data ensemble on 7 datasets, containing 4 regression and 3 classification tasks. See Appendix D.1 for details on the datasets. We use the synthetic data generators DDPM (Kotelnikov et al., 2023) and synthpop (Nowok et al., 2016), which we selected after a preliminary experiment described in Appendix E.1. We only plot the results from synthpop in the main text to save space, and defer the results of DDPM to Figure S2 in the Appendix. We generate 32 synthetic datasets, of which between 1 and 32 are used to train the ensemble. The results are averaged over 3 runs with different train-test splits.
@@
On the regression datasets, our error metric is MSE, which is the subject of Theorem 2.2. The results in Figure 1(a) show that a larger number of synthetic datasets generally decreases MSE. The decrease is especially clear with downstream algorithms that have a high variance like decision trees and 1-NN. Low-variance algorithms like ridge regression have very little if any decrease from multiple synthetic datasets. This is consistent with Theorem 2.2, where the number of synthetic datasets only affects the variance-related terms.
@@
On the classification datasets, we consider 4 error metrics. Brier score (Brier, 1950) is MSE of the class probability predictions, so Theorem 2.2 applies to it. Cross entropy is a Bregman divergence, so Theorem C.2 from Appendix C applies to it. We also included accuracy and area under the ROC curve (AUC) even though our theory does not apply to them, as they are common and interpretable error metrics, so it is interesting to see how multiple synthetic datasets affect them. We use their complements in the plots, so that lower is better for all plotted metrics. We only present the Brier score results in the main text in Figure 1(b), and defer the rest to Figures S3, S4 and S5 in Appendix E.2.
@@
Because Theorem C.2 only applies to cross entropy when averaging log probabilities instead of probabilities, we compare both ways of averaging in the Appendix E.2, but only include probability averaging in the main text.
@@
The results on the classification datasets in Figure 1(b) are similar to the regression experiment. A larger number of synthetic datasets generally decreases the score, especially for the high-variance models.
@@
In Appendix E.4, we estimate the MV and SDV terms of the decomposition in Theorem 2.2. These confirm our claims on the model variances: decision trees and 1-NN have a high variance, while linear, ridge and logistic regression have a low variance.
(a) ACS 2018
(b) German credit
Figure 2: MSE or Brier score prediction on two datasets. The predictions are very accurate on these datasets. The solid lines for DDPM and synthpop (SP-P) show the same error MSE or Brier score as Figure 1, while the dashed lines show predicted MSE or Brier score. 1-NN and 5-NN are nearest neighbours with 1 or 5 neighbours. We omitted downstream algorithms with uninteresting flat curves from panel (a). See Figure S6 in the Appendix for the full panel (a), and Figures S7 and S8 for the other datasets. Tables S11 and S12 in the Appendix contain the numbers from the plots. The results are averaged over 3 repeats with different train-test splits. The error bands are 95% confidence intervals formed by bootstrapping over the repeats.
3.2 Predicting performance from Two Synthetic Datasets
@@
Next, we evaluate the predictions our rule of thumb from Section 2.3 makes. To recap, our rule of thumb predicts that the maximal benefit from multiple synthetic datasets is $2(\mathrm{MSE}_{1}-\mathrm{MSE}_{2})$, and $m$ synthetic datasets achieve a $1-\frac{1}{m}$ fraction of this benefit, as shown in (9).
@@
Figure 2 contains the results for the ACS 2018 and German credit datasets, and Figures S6 to S8 in the Appendix contain the results for the other datasets. The predictions are very accurate on ACS 2018 and German credit, and reasonably accurate on the other datasets. The variance of the prediction depends heavily on the variance of the errors computed from the test data.
We also evaluated the rule on random forests without synthetic data, as it also applies to them. In this setting, the number of trees in the random forest is analogous to the number of synthetic datasets. We use the same datasets as in the previous experiments and use the same train-test splits of the real data. The results are in Figure S9 in the Appendix. The prediction is accurate when the test error is accurate, but can have high variance.
3.3 Differentially Private Synthetic Data
Figure 3: Brier score of the ensemble of downstream predictors with varying numbers of synthetic datasets $m$, generated with the DP methods AIM or NAPSU-MQ from the Adult dataset with a reduced set of features. Increasing the number of datasets generally decreases the score, even with AIM which splits the privacy budget between $m$ synthetic datasets. The privacy parameters are $\epsilon=1.5$, $\delta=n^{-2}\approx 4.7\cdot 10^{-7}$. The predictors are the same as in Figure 1. The black lines show the loss of the best non-DP downstream predictor trained on real data. The results are averaged over 3 repeats. The error bars are 95% confidence intervals formed by bootstrapping over the repeats. Table S13 contains the numbers from the plots, and Figure S10 contains plots of the other error metrics.
In this experiment, we evaluate the performance of the generative ensemble in the setting of Theorem 2.4, where $m$ synthetic datasets are generated from a single noisy summary $\tilde{s}$ of the real data. We compare with splitting the privacy budget between $m$ synthetic datasets.
The algorithms we use are AIM (McKenna et al., 2022), which needs to split the privacy budget between the $m$ synthetic datasets, and NAPSU-MQ (Räisä et al., 2023b), which uses a single noisy summary. The dataset is the Adult dataset with a reduced set of columns, which is needed to keep the runtime of NAPSU-MQ reasonable. The downstream task is classification, so we use the same 4 metrics, and both probability and log-probability averaging, as in the non-DP classification experiment. We use fairly strict privacy parameters of $\epsilon=1.5$, $\delta=n^{-2}\approx 4.7\cdot 10^{-7}$. See Appendix D for the full details of the setup.
The results for Brier score are in Figure 3, and the results for all metrics are in Figure S10 in the Appendix. Increasing $m$ almost always improves the results, which is expected for NAPSU-MQ due to Theorem 2.4, but somewhat surprising for AIM, which splits the privacy budget. Once again, the high-variance predictors 1-NN, 5-NN and decision trees benefit the most from multiple synthetic datasets. AIM always outperforms NAPSU-MQ, which is likely a result of AIM being able to handle a more complicated set of input queries than NAPSU-MQ.
@@ 4 Discussion
Our main theory assumes that the synthetic datasets are generated i.i.d. given either the real data or a DP summary. We generalise the bias-variance decompositions to non-i.i.d. settings in Appendix B. The implications of the decompositions, like the rule-of-thumb on the number of synthetic datasets, may not apply if i.i.d. assumption is removed.
Broader Impacts
@@ Conclusion
We derived bias-variance decompositions for using synthetic data in several cases: for MSE or Brier score with i.i.d. synthetic datasets given the real data (Section 2.2) and MSE with i.i.d. synthetic datasets given a DP summary of the real data (Section 2.4). We generalised these decompositions to non-i.i.d. synthetic datasets (Appendix B) and Bregman divergences (Appendix C). The decompositions make actionable predictions, such as yielding a simple rule of thumb that can be used to select the number of synthetic datasets. We empirically examined the performance of generative ensembles on several real datasets and downstream predictors, and found that the predictions of the theory generally hold in practice (Section 3). These findings significantly increase the theoretical understanding of generative ensembles, which is very limited in prior literature.
@@ Appendix A Missing Proofs
$\operatorname*{\mathbb{E}}_{D_{r},D_{s}^{1:m}}[\hat{g}(x;D_{s}^{1:m})]=\operatorname*{\mathbb{E}}_{D_{r}}\operatorname*{\mathbb{E}}_{D_{s}^{1:m}|D_{r}}\left[\frac{1}{m}\sum_{i=1}^{m}g(x,D_{s}^{i})\right]=\operatorname*{\mathbb{E}}_{D_{r}}\operatorname*{\mathbb{E}}_{D_{s}|D_{r}}[g(x,D_{s})]=\operatorname*{\mathbb{E}}_{D_{r},D_{s}}[g(x,D_{s})],$ (14)
@@
Appendix B Non-i.i.d. Synthetic Data
Here, we consider the case of non-i.i.d synthetic datasets, and allow each synthetic dataset to have a different predictor $g_{i}$. We get a similar decomposition as in Theorem 2.2, but the terms of the decomposition are now averages over the possibly different synthetic data distributions, and there is an additional covariance term. {theorem}[] Let the parameters for $m$ generators $\theta_{i}\sim p(\theta_{i}|D_{r})$, $i=1,\dotsc,m$, be potentially non-i.i.d. given the real data $D_{r}$. Let the synthetic datasets be $D_{s}^{i}\sim p(D_{s}^{i}|\theta_{i})$, and let $\hat{g}(x;D_{s}^{1:m})=\frac{1}{m}\sum_{i=1}^{m}g_{i}(x;D_{s}^{i})$. Then
$\mathrm{MSE}=\frac{1}{m}\overline{\mathrm{MV}}+\frac{1}{m}\overline{\mathrm{SDV}}+\mathrm{COV}+\overline{\mathrm{RDV}}+(\overline{\mathrm{SDB}}+\overline{\mathrm{MB}})^{2}+\operatorname*{Var}_{y}[y],$ (27)
where
$\overline{\mathrm{MV}}$ $=\frac{1}{m}\sum_{i=1}^{m}\operatorname*{\mathbb{E}}_{D_{r},\theta_{i}}\operatorname*{Var}_{D_{s}^{i}|\theta_{i}}[g_{i}(x;D_{s}^{i})]$ (29)
$\overline{\mathrm{SDV}}$ $=\frac{1}{m}\sum_{i=1}^{m}\operatorname*{\mathbb{E}}_{D_{r}}\operatorname*{Var}_{\theta_{i}|D_{r}}\operatorname*{\mathbb{E}}_{D_{s}^{i}|\theta_{i}}[g_{i}(x;D_{s}^{i})]$ (30)
$\mathrm{COV}$ $=\frac{1}{m^{2}}\sum_{i\neq j}\operatorname*{\mathbb{E}}_{D_{r}}\left[\operatorname*{Cov}_{D_{s}^{i},D_{s}^{j}|D_{r}}[g_{i}(x;D_{s}^{i}),g_{j}(x;D_{s}^{j})]\right]$ (31)
$\overline{\mathrm{RDV}}$ $=\operatorname*{Var}_{D_{r}}\left[\frac{1}{m}\sum_{i=1}^{m}\operatorname*{\mathbb{E}}_{D_{s}^{i}|D_{r}}[g_{i}(x;D_{s}^{i})]\right]$ (32)
$\overline{\mathrm{SDB}}$ $=\frac{1}{m}\sum_{i=1}^{m}\operatorname*{\mathbb{E}}_{D_{r}}\left[f(x)-\operatorname*{\mathbb{E}}_{\theta_{i}|D_{r}}[f_{\theta_{i}}(x)]\right]$ (33)
$\overline{\mathrm{MB}}$ $=\frac{1}{m}\sum_{i=1}^{m}\operatorname*{\mathbb{E}}_{D_{r},\theta_{i}}\left[f_{\theta_{i}}(x)-\operatorname*{\mathbb{E}}_{D_{s}^{i}|\theta_{i}}[g_{i}(x,D_{s}^{i})]\right].$ (34)
$f(x)=\operatorname*{\mathbb{E}}_{y}[y]$ is the optimal predictor for real data and $f_{\theta_{i}}$ is the optimal predictor for the synthetic data generating process given parameters $\theta_{i}$. All random quantities are implicitly conditioned on $x$.
Proof.
The classical bias-variance decomposition gives
For the bias,
$\begin{split}&f(x)-\operatorname*{\mathbb{E}}_{D_{r},D_{s}^{1:m}}[\hat{g}(x;D_{s}^{1:m})]\\ &=f(x)-\operatorname*{\mathbb{E}}_{D_{r}}\operatorname*{\mathbb{E}}_{D_{s}^{1:m}|D_{r}}\left[\frac{1}{m}\sum_{i=1}^{m}g_{i}(x,D_{s}^{i})\right]\\ &=\operatorname*{\mathbb{E}}_{D_{r}}\operatorname*{\mathbb{E}}_{D_{s}^{1:m}|D_{r}}\left[\frac{1}{m}\sum_{i=1}^{m}(f(x)-g_{i}(x,D_{s}^{i}))\right]\\ &=\operatorname*{\mathbb{E}}_{D_{r}}\operatorname*{\mathbb{E}}_{D_{s}^{1:m}|D_{r}}\left[\frac{1}{m}\sum_{i=1}^{m}(f(x)-f_{\theta_{i}}(x)+f_{\theta_{i}}(x)-g_{i}(x,D_{s}^{i}))\right]\\ &=\operatorname*{\mathbb{E}}_{D_{r}}\operatorname*{\mathbb{E}}_{\theta_{1:m}|D_{r}}\left[\frac{1}{m}\sum_{i=1}^{m}(f(x)-f_{\theta_{i}}(x))\right]+\operatorname*{\mathbb{E}}_{D_{r}}\operatorname*{\mathbb{E}}_{D_{s}^{1:m}|D_{r}}\left[\frac{1}{m}\sum_{i=1}^{m}(f_{\theta_{i}}(x)-g_{i}(x,D_{s}^{i}))\right]\\ &=\operatorname*{\mathbb{E}}_{D_{r}}\left[\frac{1}{m}\sum_{i=1}^{m}(f(x)-\operatorname*{\mathbb{E}}_{\theta_{i}|D_{r}}[f_{\theta_{i}}(x)])\right]+\operatorname*{\mathbb{E}}_{D_{r},\theta_{1:m}}\left[\frac{1}{m}\sum_{i=1}^{m}(f_{\theta_{i}}(x)-\operatorname*{\mathbb{E}}_{D_{s}^{i}|\theta_{i}}[g_{i}(x,D_{s}^{i})])\right]\\ &=\frac{1}{m}\sum_{i=1}^{m}\operatorname*{\mathbb{E}}_{D_{r}}\left[f(x)-\operatorname*{\mathbb{E}}_{\theta_{i}|D_{r}}[f_{\theta_{i}}(x)]\right]+\frac{1}{m}\sum_{i=1}^{m}\operatorname*{\mathbb{E}}_{D_{r},\theta_{i}}\left[f_{\theta_{i}}(x)-\operatorname*{\mathbb{E}}_{D_{s}^{i}|\theta_{i}}[g_{i}(x,D_{s}^{i})]\right].\end{split}$ (36)
For the variance,
$\begin{split}\operatorname*{Var}_{D_{r},D_{s}^{1:m}}[\hat{g}(x;D_{s}^{1:m})]&=\operatorname*{\mathbb{E}}_{D_{r}}\operatorname*{Var}_{D_{s}^{1:m}|D_{r}}\left[\frac{1}{m}\sum_{i=1}^{m}g_{i}(x;D_{s}^{i})\right]+\operatorname*{Var}_{D_{r}}\operatorname*{\mathbb{E}}_{D_{s}^{1:m}|D_{r}}\left[\frac{1}{m}\sum_{i=1}^{m}g_{i}(x;D_{s}^{i})\right]\\ &=\frac{1}{m^{2}}\operatorname*{\mathbb{E}}_{D_{r}}\operatorname*{Var}_{D_{s}^{1:m}|D_{r}}\left[\sum_{i=1}^{m}g_{i}(x;D_{s}^{i})\right]+\operatorname*{Var}_{D_{r}}\left[\frac{1}{m}\sum_{i=1}^{m}\operatorname*{\mathbb{E}}_{D_{s}^{i}|D_{r}}[g_{i}(x;D_{s}^{i})]\right],\end{split}$ (37)
$\begin{split}\operatorname*{Var}_{D_{s}^{1:m}|D_{r}}\left[\sum_{i=1}^{m}g_{i}(x;D_{s}^{i})\right]&=\sum_{i=1}^{m}\operatorname*{Var}_{D_{s}^{i}|D_{r}}[g_{i}(x;D_{s}^{i})]+\sum_{i\neq j}\operatorname*{Cov}_{D_{s}^{i},D_{s}^{j}|D_{r}}[g_{i}(x;D_{s}^{i}),g_{j}(x;D_{s}^{j})]\end{split}$ (38)
and
$\operatorname*{Var}_{D_{s}^{i}|D_{r}}[g_{i}(x;D_{s}^{i})]=\operatorname*{\mathbb{E}}_{\theta_{i}|D_{r}}\operatorname*{Var}_{D_{s}^{i}|\theta_{i}}[g_{i}(x;D_{s}^{i})]+\operatorname*{Var}_{\theta_{i}|D_{r}}\operatorname*{\mathbb{E}}_{D_{s}^{i}|\theta}[g_{i}(x,D_{s}^{i})].$ (39)
If the synthetic datasets are identically distributed, but not necessarily independent, and the predictors are identical, Theorem B simplifies to
$\mathrm{MSE}=\frac{1}{m}\mathrm{MV}+\frac{1}{m}\mathrm{SDV}+\left(1-\frac{1}{m}\right)\mathrm{COV}+\mathrm{RDV}+(\mathrm{SDB}+\mathrm{MB})^{2}+\operatorname*{Var}_{y}[y],$ (40)
where
$\mathrm{COV}$ $=\operatorname*{\mathbb{E}}_{D_{r}}\left[\operatorname*{Cov}_{D_{s}^{1},D_{s}^{2}|D_{r}}[g(x;D_{s}^{1}),g(x;D_{s}^{2})]\right]$ (44)
which is Theorem 2.2 with the additional covariance term.
In the noisy summary case, we get an analogue of Theorem 2.4. {theorem}[] Let the parameters for $m$ generators $\theta_{i}\sim p(\theta_{i}|\tilde{s})$, $i=1,\dotsc,m$, be potentially non-i.i.d. given a DP summary $\tilde{s}$ the real data $D_{r}$, let the synthetic datasets be $D_{s}^{i}\sim p(D_{s}^{i}|\theta_{i})$, and let $\hat{g}(x;D_{s}^{1:m})=\frac{1}{m}\sum_{i=1}^{m}g_{i}(x;D_{s}^{i})$. Then
$\mathrm{MSE}=\frac{1}{m}\overline{\mathrm{MV}}+\frac{1}{m}\overline{\mathrm{SDV}}+\mathrm{COV}+\overline{\mathrm{RDV}}+\overline{\mathrm{DPVAR}}+(\overline{\mathrm{SDB}}+\overline{\mathrm{MB}})^{2}+\operatorname*{Var}_{y}[y],$ (48)
where
$\overline{\mathrm{MV}}$ $=\frac{1}{m}\sum_{i=1}^{m}\operatorname*{\mathbb{E}}_{D_{r},\tilde{s},\theta_{i}}\operatorname*{Var}_{D_{s}^{i}|\theta_{i}}[g_{i}(x;D_{s}^{i})]$ (50)
$\overline{\mathrm{SDV}}$ $=\frac{1}{m}\sum_{i=1}^{m}\operatorname*{\mathbb{E}}_{D_{r},\tilde{s}}\operatorname*{Var}_{\theta_{i}|\tilde{s}}\operatorname*{\mathbb{E}}_{D_{s}^{i}|\theta_{i}}[g_{i}(x;D_{s}^{i})]$ (51)
$\mathrm{COV}$ $=\frac{1}{m^{2}}\sum_{i\neq j}\operatorname*{\mathbb{E}}_{D_{r},\tilde{s}}\left[\operatorname*{Cov}_{D_{s}^{i},D_{s}^{j}|\tilde{s}}[g_{i}(x;D_{s}^{i}),g_{j}(x;D_{s}^{j})]\right]$ (52)
$\overline{\mathrm{RDV}}$ $=\operatorname*{Var}_{D_{r}}\left[\frac{1}{m}\sum_{i=1}^{m}\operatorname*{\mathbb{E}}_{D_{s}^{i}|D_{r}}[g_{i}(x;D_{s}^{i})]\right]$ (53)
$\overline{\mathrm{DPVAR}}$ $=\operatorname*{\mathbb{E}}_{D_{r}}\operatorname*{Var}_{\tilde{s}|D_{r}}\left[\frac{1}{m}\sum_{i=1}^{m}\operatorname*{\mathbb{E}}_{D_{s}^{i}|\tilde{s}}[g_{i}(x;D_{s}^{i})]\right]$ (54)
$\overline{\mathrm{SDB}}$ $=\frac{1}{m}\sum_{i=1}^{m}\operatorname*{\mathbb{E}}_{D_{r},\tilde{s}}\left[f(x)-\operatorname*{\mathbb{E}}_{\theta_{i}|\tilde{s}}[f_{\theta_{i}}(x)]\right]$ (55)
$\overline{\mathrm{MB}}$ $=\frac{1}{m}\sum_{i=1}^{m}\operatorname*{\mathbb{E}}_{D_{r},\tilde{s},\theta_{i}}\left[f_{\theta_{i}}(x)-\operatorname*{\mathbb{E}}_{D_{s}^{i}|\theta_{i}}[g_{i}(x,D_{s}^{i})]\right].$ (56)
$f(x)=\operatorname*{\mathbb{E}}_{y}[y]$ is the optimal predictor for real data and $f_{\theta_{i}}$ is the optimal predictor for the synthetic data generating process given parameters $\theta_{i}$. All random quantities are implicitly conditioned on $x$.
Proof.
Using $\tilde{s}$ in place of $D_{r}$ in Theorem B gives
$\mathrm{MSE}=\frac{1}{m}\overline{\mathrm{MV}}+\frac{1}{m}\overline{\mathrm{SDV}}+\mathrm{COV}+\operatorname*{Var}_{D_{r},\tilde{s}}\left[\frac{1}{m}\sum_{i=1}^{m}\operatorname*{\mathbb{E}}_{D_{s}^{i}|\tilde{s}}[g_{i}(x;D_{s}^{i})]\right]+(\overline{\mathrm{SDB}}+\overline{\mathrm{MB}})^{2}+\operatorname*{Var}_{y}[y].$ (57)
The variance over $D_{r}$ and $\tilde{s}$ can be decomposed
$\begin{split}&\operatorname*{Var}_{D_{r},\tilde{s}}\left[\frac{1}{m}\sum_{i=1}^{m}\operatorname*{\mathbb{E}}_{D_{s}^{i}|\tilde{s}}[g_{i}(x;D_{s}^{i})]\right]\\ &=\operatorname*{\mathbb{E}}_{D_{r}}\operatorname*{Var}_{\tilde{s}|D_{r}}\left[\frac{1}{m}\sum_{i=1}^{m}\operatorname*{\mathbb{E}}_{D_{s}^{i}|\tilde{s}}[g_{i}(x;D_{s}^{i})]\right]+\operatorname*{Var}_{D_{r}}\operatorname*{\mathbb{E}}_{\tilde{s}|D_{r}}\left[\frac{1}{m}\sum_{i=1}^{m}\operatorname*{\mathbb{E}}_{D_{s}^{i}|\tilde{s}}[g_{i}(x;D_{s}^{i})]\right]\\ &=\operatorname*{\mathbb{E}}_{D_{r}}\operatorname*{Var}_{\tilde{s}|D_{r}}\left[\frac{1}{m}\sum_{i=1}^{m}\operatorname*{\mathbb{E}}_{D_{s}^{i}|\tilde{s}}[g_{i}(x;D_{s}^{i})]\right]+\operatorname*{Var}_{D_{r}}\left[\frac{1}{m}\sum_{i=1}^{m}\operatorname*{\mathbb{E}}_{D_{s}^{i}|D_{r}}[g_{i}(x;D_{s}^{i})]\right].\end{split}$ (58)
Appendix C Bias-Variance Decomposition for Bregman Divergences
C.1 Background: Bregman Divergences